### v1

# Online Budget-Feasible Mechanism Design with Predictions

Abstract
Augmenting the input of algorithms with predictions is an algorithm design paradigm that suggests leveraging a (possibly erroneous) prediction to improve worst-case performance guarantees when the prediction is perfect (consistency), while also providing a performance guarantee when the prediction fails (robustness). Recently, xu2022mechanism and agrawal2022learning proposed to consider settings with strategic agents under this framework. In this paper, we initiate the study of budget-feasible mechanism design with predictions. These mechanisms model a procurement auction scenario in which an auctioneer (buyer) with a strict budget constraint seeks to purchase goods or services from a set of strategic agents, so as to maximize her own valuation function. We focus on the online version of the problem where the arrival order of agents is random. We design mechanisms that are truthful, budget-feasible, and achieve a significantly improved competitive ratio for both monotone and non-monotone submodular valuation functions compared to their state-of-the-art counterparts without predictions. Our results assume access to a prediction for the value of the optimal solution to the offline problem. We complement our positive results by showing that for the offline version of the problem, access to predictions is mostly ineffective in improving approximation guarantees.
1 Introduction
Our work evolves around the design of auction mechanisms for procurement auctions. Procurement auctions, also known as reverse auctions, are mechanisms for assigning a set of projects or tasks to a set of candidate providers. In the more traditional applications of reverse auctions, the auctioneer may be a governmental organization (e.g., for assigning a public project to a construction firm), or a company interested in outsourcing tasks, such as the supply of logistics or retail services; see CSS06. In recent years, the explosion and monetization of the Internet, have created even further applications and platforms where even a single individual can conduct online assignments/auctions. Two such prominent families of examples include platforms for online labour markets, such as Upwork or Freelancer, and platforms for crowdsourcing systems, where the posted tasks can range from very elementary assignments as image labeling, all the way to more specialized services like translating documents. Participatory crowdsensing also falls under this framework, where some entities are interested in collecting sensing information from smartphones or other devices.
The emergence of such online markets, has caught the increasing attention of the research community, which has focused on properly formulating and addressing relevant research questions. One of the popular models in the literature for capturing procurement auctions is that of budget-feasible mechanism design, originally proposed in the seminal paper of singer2010budget. This model concerns a scenario where an auctioneer with a strict budget constraint seeks to purchase goods or services from a set of strategic agents, who may misreport their cost to their advantage for obtaining higher payments. Under these considerations, a natural goal for the auctioneer is to come up with a truthful mechanism for hiring a subset of the agents that maximizes the procured value (measured by the valuation function of the auctioneer), and such that the total payments to the agents respect the budget limitation. Given that even the non-strategic versions of such budget-constrained problems tend to be NP-hard, the main focus is on providing budget-feasible mechanisms that achieve approximation guarantees on the auctioneer’s optimal potential value. This has led to a steady stream of works which we outline in our related work section.
The focus of our work is to study budget-feasible mechanisms under the rather recently introduced framework of learning-augmented mechanism design. This forms a new paradigm in designing algorithms, where the input is augmented by a prediction on some relevant parameter of the problem, possibly obtained by past data. The main goal is to leverage a (possibly erroneous) prediction to improve worst-case performance guarantees when the prediction is accurate (referred to as consistency), while also providing a performance guarantee when it is not (referred to as robustness), without knowing the prediction accuracy in advance. Recently, xu2022mechanism and agrawal2022learning proposed to consider settings with strategic agents through this lens. Motivated by these two initial attempts, there has been a fast growing interest and a series of other works, that studied the design of truthful mechanisms with predictions in a variety of settings, leading mostly to positive results (i.e., improved approximation guarantees under correct predictions while not far off from the best known guarantee otherwise).
Our Contribution.
We initiate the study of budget-feasible mechanism design with predictions. We focus on the online version of the problem in the random order model, where the agents arrive in a uniformly random order, and under a submodular valuation function for the auctioneer. Further, the input of our mechanisms is augmented by a prediction for the value of the optimal solution to the offline problem. Our main results can be summarized as follows:
  • In Section 3 we propose a family of universally truthful, budget-feasible mechanisms (LABEL:mechanism:mono-sm-conv; parameterized by an error tolerance parameter $\tau$) for monotone submodular objectives which attain $O(1)$ consistency and robustness. For $\tau$ close to $1$, the consistency, i.e., the approximation guarantee when the prediction is perfect, can be as low as $6$, which is a significant improvement even versus the best known approximation guarantee of $54.4$ for Submodular Knapsack Secretary, i.e., the non-strategic version of the problem, without predictions (due to feldman11).
  • For smaller values of $\tau$, the robustness of LABEL:mechanism:mono-sm-conv in Section 3, i.e., the approximation guarantee when the prediction is arbitrarily bad, can be as low as $146$. This also implies a $146$-approximation mechanism without predictions, improving the previously known approximation of $1710$ by amanatidis2022budget by a factor greater than $11$!
  • In LABEL:sec:mono-sm-sampling, we present a different mechanism (LABEL:mechanism:mono-sm2) where the prediction implicitly determins the length of the initial agent-sampling window (although here we have fixed the latter for the sake of presentation). Compared to LABEL:mechanism:mono-sm-conv, this approach is more conservative, as LABEL:mechanism:mono-sm2 is, in a sense, less sensitive to (very) erroneous predictions. In particular, we show that our mechanism achieves a consistency-robustness tradeoff of $95$ and $280$.
  • We obtain analogous results for non-monotone submodular objectives, albeit with some loss on approximation, via LABEL:mechanism:nonmono-sm-conv and LABEL:mechanism:non-mono-sm2. Specifically, for $\tau$ close to $1$, the consistency of LABEL:mechanism:nonmono-sm-conv can be as low as $19$, whereas, for smaller values of $\tau$, its robustness can be as low as $445$. We emphasize once more that, even when one completely ignores the prediction, our analysis leads to a large improvement on the current state-of-the-art approximation due to amanatidis2022budget, which is $1710$ even for Submodular Knapsack Secretary (the non-strategic version of the problem). Finally, we show that the result of the analogue of LABEL:mechanism:mono-sm2 for non-monotone submodular valuations (LABEL:mechanism:non-mono-sm2) achieves a consistency-robustness tradeoff of $228$ and $818$.
  • We complement our positive results by showing in LABEL:sec:lower-bound that for the offline version, access to predictions is mostly ineffective in improving approximation guarantees. In particular, we show that no randomized, universally truthful, budget-feasible mechanism with bounded robustness can achieve a consistency guarantee better than the lower bound of $2$ that is known to hold for additive valuations and randomized mechanisms (chen2011approximability).
All our mechanisms run in polynomial time assuming access to a value oracle for the objective function, which is standard in submodular optimization. From a technical point of view, we design posted price auction mechanisms, the prices of which are determined by exploiting the information from the prediction on the optimal value and the outcome of an initial sampling phase. Overall, our results shed light on the power but also on the limitations of the learning-augmented paradigm for budget-feasible mechanisms.
Related Work.
The design of truthful budget-feasible mechanisms was initiated by singer2010budget for additive and monotone submodular valuation functions, and has sparked a rich line of work. For additive valuation functions a best-possible $2$-approximation randomized mechanism was given by gravin2020optimal, whereas it is also known that the best possible factor of any deterministic mechanism is between $1+\sqrt{2}$ and $3$ (due to chen2011approximability and gravin2020optimal). For monotone submodular valuation functions, through a series of improvements (see e.g., chen2011approximability; jalaly2021simple; balkanski2022deterministic), the current state of the art is $4.45$ for deterministic and $4.08$ for randomized mechanisms due to han2024triple who follow the clock auction paradigm. Without the monotonicity assumption, obtaining any $O(1)$ approximation seemed to be more challenging (see e.g., amanatidis2022budget), but here as well the recent improvements of balkanski2022deterministic; huang2023randomized have led to a relatively small approximation ratio, namely $11.67$, again by han2024triple.
There have also been works focusing on richer valuation classes like XOS and subadditive functions (see e.g. bei2017worst; amanatidis2017budget; neogi_et_al:LIPIcs.ITCS.2024.84), exploring beyond-the-worst-case scenarios (rubinstein23), or using large market assumptions, see AnariGN18; jalaly2021simple. We refer the reader to the recent survey of liu24 for an overview of various settings (e.g., environments with combinatorial constraints and more general single-parameter domains).
A recent direction in theoretical computer science is the area of “beyond the worst-case analysis” (see, e.g., roughgarden2021beyond). One prominent approach within this framework is algorithms with predictions, which originated in the field of online algorithms (see, e.g., lykouris2021competitive) and utilizes predictions, e.g., provided by an ML algorithm trained on historical data, to overcome worst case lower bounds. The idea has been also applied in strategic settings, where predictions are used to simulate (an aggregation of) the agents’ private information, starting with the work of agrawal2022learning and xu2022mechanism. Areas of application in mechanism design include auction-related environments (see e.g., prasad23; balkanski23; christodoulou24; caragiannis24; balkanski23scheduling; gkatzelis25), online mechanisms (see e.g., balkanski23; lu24), mechanisms without money (see e.g. balkanski24randomized; colini24; barak24; cohen24) and computational social choice (see e.g. filos2025). For an almost exhaustive list of papers in the area, see the online repository of algorithms_with_predictions.
The online budget-feasible mechanism design problem we study was introduced by badanidiyuru2012learning and is a generalization to the strategic setting of the Submodular Knapsack Secretary problem (see, e.g., babaioff17). The latter has also been considered by bateni13 and subsequently by feldman11 and KesselheimT17. Very recently, charalampopoulos25 devised a truthful mechanism which achieves a constant approximation (for a very large constant) for posted pricing and monotone submodular valuations. The state-of-the-art approximation for monotone or non-monotone submodular valuation functions for the strategic version we study here is due to amanatidis2022budget, who achieve a competitive ratio of $1710$. Note that the Secretary problem (additive objective and cardinality constraint) has been studied in environments with predictions (see, e.g., antoniadis23; perchet23; fujii24; balkanski24fair), yet our work is the first to study the online budget-feasible mechanism design problem—and its algorithmic counterpart—in such augmented environments.
2 Preliminaries
We explore online mechanism design with predictions, focusing on a procurement auction setting involving a single auctioneer and multiple agents. The auctioneer has a valuation function $v:2^{N}\rightarrow\mathbb{Q}_{\geq 0}$ and a budget $B>0$. We use $N=[n]=\{1,2,\dots,n\}$ to denote the set of $n$ agents. Each agent $i$ has a private cost $c_{i}>0$, namely the cost of getting hired by the auctioneer. For $S\subseteq N$, $v(S)$ represents the value the auctioneer derives from selecting that set; for singletons, we write $v(i)$ instead of $v(\{i\})$. The algorithmic goal in all the problems we study is to select a set $S$ that maximizes $v(S)$ subject to the constraint $\sum_{i\in S}c_{i}\leq B$. We assume oracle access to $v$ via value queries, implying the existence of a polynomial-time value oracle that returns $v(S)$ when queried with a set $S$. A function $v$ is non-decreasing (referred to as monotone herein), if $v(S)\leq v(T)$ for any $S\subseteq T\subseteq N$. We consider the cases of monotone and non-monotone, normalised (i.e., $v(\emptyset)=0$), non-negative submodular valuation functions. Since marginal values are extensively used, we adopt the shortcut $v(i|S)$ for the marginal value of agent $i$ with respect to the set $S$, i.e., $v(i|S):=v(S\cup\{i\})-v(S)$.
Definition 2.1.
A function $v$, defined on $2^{N}$ for some set $N$, is submodular if $v(i|S)\geq v(i|T)$ for all $S\subseteq T$ and $i\notin T$.
Note that when $v(i|S)=v(i|\emptyset)$, for all $i\in N$ and all $S\subseteq N$, the valuation function $v$ is additive. The following known result will be useful in our analysis in Section 3.
Theorem 2.2 (NemhauserWF78).
A function $v$ is submodular if and only if, for all $S,T\subseteq N$
$v(T)\leq v(S)+\textstyle\sum_{i\in T\setminus S}v(i|S)+\textstyle\sum_{i\in S\setminus T}v(i|(S\cup T)\setminus\{i\}).$
An important element in our problem is the online arrival of agents. We adopt the random order (RO) model; see also the discussion in the beginning of Section 3. In this model an adversary first chooses the costs of all the agents in $N$. There are $n$ different time slots $t_{1},\dots,t_{n}$, with a single agent appearing in each slot according to a uniformly random permutation $\pi$, i.e., the input sequence to an online algorithm is ${\pi(1)},{\pi(2)},\dots,{\pi(n)}$. Such an algorithm has to perform its irrevocable actions for agent ${\pi(i)}$ before seeing agent ${\pi(i+1)}$. The length $n$ of the input sequence may be known in advance, and we assume so in our setting. For the RO model, given the output $A$ of an algorithm and the value of an optimal solution $OPT$, we define the competitive-ratio (for maximization problems) to be $\frac{OPT}{\mathbb{E}[v(A)]}$ on the adversarily-chosen (worst case) set of inputs. The expected value is over all permutations of the input.
Mechanism Design. An offline mechanism $M=(A,p)$ consists of an allocation function $A:\mathbb{R}_{\geq 0}^{n}\rightarrow 2^{[n]}$ and a payment rule $p:\mathbb{R}_{\geq 0}^{n}\rightarrow\mathbb{R}_{\geq 0}^{n}$. Given a profile of cost declarations $\mathbf{b}\in[0,B]^{n}$, where each $b_{i}$ is the cost reported by agent $i$, the allocation function selects a set of agents $A(\mathbf{b})\subseteq N$. The payment rule returns a profile of payments $\mathbf{p}(\mathbf{b})=(p_{1},\dots,p_{n})$, where $p_{i}(\mathbf{b})\geq 0$ is the payment to agent $i\in N$. Therefore, the utility of agent $i\in N$ is $u_{i}(\mathbf{b})=p_{i}(\mathbf{b})-c_{i}$ if $i\in A(\mathbf{b})$ and $u_{i}(\mathbf{b})=0$ otherwise, with the understanding that agents not in $A(\mathbf{b})$ are not paid.
An online mechanism, on the other hand, decides the allocation and payment of an agent at every time step without having seen the entire input. That is, given a permutation $\pi$ on $N$, the mechanism decides about agent $\pi(t)$ at time step $t\in[n]$, as $\pi(t)$ appears and declares her cost, i.e., at the time the mechanism has only seen $\mathbf{b}_{t}^{\pi}=(b_{\pi(1)},\dots,b_{\pi(t)})\in[0,B]^{t}$. The mechanism, thus, maintains the pair $(A_{t}(\mathbf{b}_{t}^{\pi}),p_{t}(\mathbf{b}_{t}^{\pi}))$ of currently selected agents and their payments, which is consistent with the decisions being irrevocable.
For a deterministic mechanism $M=(A,p)$ we require that, for every fixed arrival order, any true cost profile $\mathbf{c}$ and any declared cost profile $\mathbf{b}$, the mechanism is:
  • budget-feasible; the payments to the agents do not exceed the auctioneer’s budget, i.e., $\sum_{i=1}^{n}p_{i}(\mathbf{b})\leq B.$
  • individually rational; no agent can be hired for less than they asked for, i.e., $p_{i}(\mathbf{b})\geq b_{i}$, if $i\in A(\mathbf{b})$
  • truthful; no agent $i\in N$ has an incentive to misreport their true cost $c_{i}$, i.e., $u_{i}(c_{i},\mathbf{b}_{-i})\geq u_{i}(\mathbf{b})$.
A randomized mechanism $M=(A,p)$ can be thought of as a probability distribution over deterministic mechanisms for every fixed arrival order. In this work, we require all our randomized mechanisms to be probability distributions over budget-feasible, individually rational, and truthful deterministic mechanisms. Randomized mechanisms with the latter property are called universally truthful. [^1] This is stronger than truthfulness in expectation, which requires that truth-telling maximizes each agent’s expected utility. Since all our mechanisms are provably universally truthful, we will consistently refer to the true cost profile $\mathbf{c}$ instead of the declared cost profile $\mathbf{b}$ when needed. Furthermore, we will omit referring explicitly to $\mathbf{c}$ if it is clear from the context.
Given the above, an instance of the problem is described by the tuple $I=(N,\mathbf{c},v,B)$. A common practice in this line of work is to measure the performance of a mechanism $M=(A,p)$, both in online variants and the offline problem, against the optimal value of the following packing problem:
$\max\ v(S)\quad\text{s.t.}\quad\textstyle\sum_{i\in S}c_{i}\leq B,\quad S\subseteq N.\vspace{-3pt}$ (1)
We use $S^{*}(\mathbf{c})\subseteq N$ to refer to an optimal solution of (1).
Predictions. We employ a beyond-worst-case analysis perspective to measure the performance of mechanisms. We assume that each instance $I=(N,\mathbf{c},v,B)$ is augmented by a (possibly erroneous) deterministic prediction $\omega$ of $v(S^{*}(\mathbf{c}))$, and that a mechanism is equipped with an error tolerance parameter $\tau\in[0,1)$, set by the mechanism designer, representing how much tolerance she has for the prediction’s inaccuracy. We denote the augmented instance as $I^{+}=(N,\allowbreak\mathbf{c},v,B,\omega)$. As is standard in the algorithms with predictions literature, we assess the performance of a mechanism with tolerance $\tau$, $M_{\tau}=(A,p)$, on augmented instances based on the following two metrics:
  • Consistency: $M_{\tau}$ is $\alpha$-consistent if $\alpha\cdot\operatorname*{\mathbb{E}}\left[v(A(\mathbf{c}))\right]\geq v(S^{*}(\mathbf{c}))$, for every instance with a perfect prediction, i.e., when $\omega=v(S^{*}(\mathbf{c}))$.
  • Robustness: $M_{\tau}$ is $\beta$-robust if $\beta\cdot\operatorname*{\mathbb{E}}\left[v(A(\mathbf{c}))\right]\geq v(S^{*}(\mathbf{c})),$ for every instance with an arbitrary prediction $\omega\in(0,v(S^{*}(\mathbf{c}))]$.
Clearly, $\alpha,\beta\geq 1$. Also, note that the expectation in each of the two benchmarks above accounts for both the randomness of $M_{\tau}$ and the randomness of the RO model. In addition to consistency and robustness, we also assess the performance of mechanisms when the prediction is “approximately” correct. To this end, we introduce an error parameter $\varepsilon$ that quantifies how far the prediction is from $v(S^{*}(\mathbf{c}))$. Formally, for an instance $I^{+}$, we define the constant $\varepsilon\in[0,1)$ to be such that $\omega=(1-\varepsilon)\cdot v(S^{*}(\mathbf{c}))$, or equivalently, $\varepsilon=1-\omega/v(S^{*}(\mathbf{c}))$.
Note that, throughout this work, we assume that $\omega$ is an underestimator of $v(S^{*})$ and, thus, the analyses presented refer to a parameterization of the prediction as $\omega=(1-\varepsilon)\cdot v(S^{*})$ with $\varepsilon\in[0,1)$. This is mainly for the sake of presentation; our results extend for $\omega=(1+\varepsilon)\cdot v(S^{*})$ with $\varepsilon\in[0,\kappa]$, for constant $\kappa$, through very similar analyses.
3 A Natural Prediction-Augmented Mechanism
In this section, we assume that the valuation function $v$ is monotone submodular. We address the more general, non-monotone case in LABEL:sec:nonmono-submodular.
We start with a short discussion about the online model we consider. Clearly, the most well-studied model for online algorithms, is the adversarial one. For our problem, that would mean that an adversary chooses the cost and value of each agent, as well as the time of their arrival. Versus such an adversary though, no meaningful approximation guarantee for consistency and robustness simultaneously is possible.
Proposition 3.1 (name=).
No online mechanism, augmented with a prediction for the value of an optimal solution, can provide both a bounded consistency and a bounded robustness guarantee in the adversarial model of arrival.
Proof.
An online mechanism augmented with a prediction for the value of the optimal solution, cannot provide both a bounded consistency and bounded robustness guarantee. To see this, consider the scenario where all the agents have the same cost equal to $B$, agent $i$ has value $v(i)=1$ and for each other agent $j\neq i$, $v(j)=\epsilon$, for some small $\epsilon>0$. Obviously, in this case the optimal solution $S^{*}$ consists only of agent $i$. Any mechanism, that tries to gain information through a sampling phase, has an unbounded consistency and robustness guarantee for such an instance, since the adversary can always select the time of arrival of agent $i$ in a way that she always belongs to the sampled set. Moreover, should a mechanism omit the sampling phase, then it should either select an agent arbitrarily at random, which again does not provide any meaningful approximation guarantee, or select an agent based on the information of the prediction, which will lead to an unbounded robustness factor, since an erroneous prediction can lead the mechanism to select an agent $j$, with $v(j)=\epsilon$. ∎
Given Proposition 3.1, we turn our attention on the RO model of arrival, defined in Section 2. The main result of this section is LABEL:mechanism:mono-sm-conv, which is a universally truthful, budget-feasible randomized mechanism with constant consistency and robustness guarantees that runs in polynomial time. Moreover, the tradeoff between consistency and robustness that the mechanism achieves is tunable via an error tolerance parameter $\tau\in[0,1)$. LABEL:mechanism:mono-sm-conv can be viewed as a convex combination between two other mechanisms, namely LABEL:mechanism:mono-sm-pred and LABEL:mechanism:mono-sm-withoutpred. The first one (LABEL:mechanism:mono-sm-pred), is a posted price mechanism that builds a budget feasible solution based solely on the information provided by the prediction. The second one (LABEL:mechanism:mono-sm-withoutpred), is also a posted price mechanism which, however, ignores the prediction. This mechanism bears similarities to the mechanism of badanidiyuru2012learning, however, our analysis is significantly tighter. The resulting mechanism (LABEL:mechanism:mono-sm-conv) randomizes between the two outcomes using a tolerance parameter $\tau\in[0,1)$.
[…]